Status Report

NASA Space Station On-Orbit Status 11 Dec 2003

By SpaceRef Editor
December 11, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 11 Dec 2003
iss

All ISS systems continue to function nominally, except as noted previously or below.

Yesterday’s re-assembly of the TVIS treadmill in the Service Module (SM) was successfully completed, and the machine is now back in the “pit”. Today the crew conducted an activation checkout, viz., inspection, video documentation (with base skirt removed) and exercise. Go-ahead for normal ops is expected on Friday (12/12). [The video and exercise data on PCMCIA card (portable computer memory card international adapter) were downlinked for engineering assessment and approval of resumption of normal TVIS operation in the contingency configuration. During the two-day IFM (in-flight maintenance) the top bearing of the roll-stabilizing gyro was found to be bad, five (instead of four) frayed wire ropes and fasteners were replaced, and the processor was reconfigured, via DIP switch, to enable yaw and pitch axis stabilization without the gyro. After approval, normal TVIS ops will involve non-motorized mode, with stabilizers but no gyro, and with the SPDs (subject positioning devices) in out-rigged config, valid until 3/1/04. An additional TVIS test with higher subject loading will be conducted in February because of the increased subject loading in March due to the crew’s heavier workout load starting 30 days prior to their return to Earth gravity.]

After connecting the UOP-DCP (utility outlet panel-to-display & control panel) bypass power cable at the Lab RWS (robotics workstation) in the morning, Mike Foale took the SSRMS (Space Station Remote Manipulator System) through the scheduled proficiency training session. [The procedure’s objective was an outside inspection with the MSS video cameras of (1) the FGB, (2) the SM portside (to check for any indication of the noise heard on 11/26, (3) the P6 truss’ 4-bar linkage shoulder bolts (to determine if the bolts have backed out), (4) the ISS zenith (top) side, and (5) finally moving to the pre-grapple position at the PDGF-1 (power & data grapple fixture #1). All steps were accomplished except the grapple (to be done at a later time). Cursory visual inspection of the exterior did not show anything unusual. The downlinked video will be carefully analyzed on the ground. After the session, Mike safed the RWS and disconnected the UOP-DCP cable.]

FE Alex Kaleri configured the Progress Rodnik tank system for liquid waste (urine) transfer from the SM’s holding tank to the empty Rodnik water tanks. The transfer, from about a dozen EDV containers to the SM tankage and thence to the Progress, proceeded during the day, driven by a compressor.

CDR Foale used the U.S. sound level meter (SLM) to take a 2-hr. acoustic survey at 41 locations in the Lab, Node, Airlock, FGB, SM and DC-1 docking compartment. [These acoustic measurements are obtained every other month. The SLM gives instantaneous noise levels and their frequency spectra, which are transferred to the MEC (medical equipment computer) laptop via an RS232 cable and later downlinked with regular CHeCS (crew health care systems) data dump or via OCA.]

At 12:45pm EST, Mike Foale conferred via S-band with key Science personnel to discuss ISS payload work. [Participants in the 15-min. teleconference included the ISS Program Scientist, the Lead Increment Scientist, the Increment Payload Manager, and the Payload Operations Manager.]

Sasha Kaleri prepared the daily IMS delta file for automatic export/import to update the database; he also performed the daily routine maintenance of the SOZh life support system (including ASU toilet facilities).

At 7:45am, both crewmembers supported two interactive TV PAO interviews of 10 min. duration each, the first with ABC Radio Network (Kenny King), and the second with Steven Young of spaceflightnow.com.

Tomorrow at 10:05am EST, Foale and Kaleri are scheduled for an interactive PAO/educational TV downlink to a Centennial of Flight Educational Event at the Wright Brothers National Memorial (National Parks Service), Kill Devil Hills, NC, with students from Kill Devil Hills and Kitty Hawk schools. The event will be live on NASA TV.

This morning, a PCS (portable computer system) laptop went down due to failure of its graphics card, leaving the crew with only two operating PCS machines, one in the Lab, the other in the SM. [To support the Robotics operations today, the Lab PCS was moved from the Cupola RWS to the Lab RWS. A third functioning machine will be cobbled together by combining the shell of the Airlock PCS, which has a bad hard drive, with the failed laptop’s hard drive. Loading (“ghosting”) of an additional PCS hard drive is scheduled for next week.]

Moscow is investigating the cause of the recent (12/9) failure of the Elektron electrolysis unit, which was down for ~10 hours. [One suspect is the possibility of an air bubble, in spite of the prior air/liquid separation procedure. The pressure jump associated with a bubble could have triggered the primary pump’s switchover to the backup pump, which then drew the bubble to itself and shut down also. Another explanation could be a computer glitch. The O2 generator was successfully restarted yesterday morning at 2:14am EST and has been running nominally ever since. There is also about 12 days’ worth of gaseous O2 stored in the Progress 12P resupply tank.]

Yesterday, the FDIR (failure detection, isolation, and recovery system) in the U.S. segment unexpectedly executed a handover of the primary internal audio controller #2 (IAC-2) to the backup IAC-1. There is no impact at this time. [IAC-1 remains active. IAC-2, currently powered off, was found to be ready to function if selected by the audio FDIR, which is enabled.]

Last Thursday’s (12/4) failure of a Microgravity Science Glovebox (MSG) circuit breaker was caused by a 6-amps current spike from the PFMI (Pore Formation & Mobility Investigation) hardware. Troubleshooting is underway. [PFMI is deactivated. Data and photographs of the thermal chamber were downlinked for analysis. An OCR (on-orbit checkout requirement) is expected for next week for pulling out the sample (leaving it bungeed inside the MSG for future troubleshooting), then unfastening the thermal chamber and rotating it to check on possible binding in its gears.]

New payload operations software (version R4) is being uplinked today for loading onto the two Payload MDMs (multiplexer/demultiplexer, computers).

Planning continues at TsUP/Moscow for a major R&R (removal & replacement) involving the BTA heat exchanger of the Russian segment’s SKV-2 air conditioner, starting next week on 12/16. [Changing the BTA heat exchanger on the SKV-2’s BK condenser unit was decided on because current inefficient performance of SKV-2 equates to lack of redundancy for the operating SKV-1. The R&R activity, performed similarly by Expedition 4’s Yuri Onufrienko and Carl Walz on SKV-1 on 4/3/2002, was originally planned for Increment 7 but was cancelled since the safety certificate for the necessary liquid Khladon-218 (Freon) venting to space was not available (it is now). The R&R will be split into four stages of one day each, with crew involvement in three of them. A crew review of the planned work is scheduled tomorrow.]

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were SW Sudan swamplands (sun glint opportunity for a mapping swath right of track–for precise documentation of the deranged stream patterns of this region on the edge of the desert where stream dynamics change drastically. Recent ISS/CEO sun glint imagery of equally flat landscapes in central Argentina reveal many patterns that available contour maps fail to show), Lagos, Nigeria (nadir pass. Lagos sits on both sides of the coastal lagoon), Lower Amazon River Basin (400-mm lens: Nadir pass over the estuary. Looking left and right for detail of rapidly changing coastlines of islands in the estuary. Islands move en masse downstream at rates that are still under study, and are of great interest as analogs for understanding rivers sediments on both Earth and Mars), Berlin, Germany (looking a touch north. Opportunity to take advantage of clear weather. Although winter lighting conditions are not optimal, 180-mm lens images can be enhanced to reveal the extent of built-up areas), Monterrey, Mexico (looking a touch left for Mexico’s second city, which has a strongly radial pattern at the foot of local mountains), and Denver, Colorado (nadir pass).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:20pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 20A. Vozdukh CO2 scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.7; ppO2 (mmHg) — 149.0; ppCO2 (mmHg) — 3.9.
  • SM Transfer Compartment: Pressure (mmHg) — 761; temperature (deg C) — 20.0.
  • FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 21.0.
  • Node: Pressure (mmHg) — 748.98; temperature (deg C) — 23.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 751.08; temperature (deg C) — 24.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 751.19; temperature (deg C) — 26.1; shell heater temp (deg C) — 24.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 22.7
  • PMA-2: Shell heater temp (deg C) — 12.0

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking).
  • SM batteries: Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting, off-line; all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries: Battery #5 is off (capacity restoration mode, ROM); battery #6 is in “Cycle Mode”; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available:3703 kg (8164 lb) as of 12/4/03 [SM (755) + FGB (2596) + Progress M (352) + Progress M-1 (0)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: 90.2 deg, pitch: -10.8 deg, roll: -3 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 6:57am EST [= epoch]):

  • Mean altitude — 371.7 km
  • Apogee — 375.7 km
  • Perigee — 367.7 km
  • Period — 92.0 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005927
  • Orbits per 24-hr. day — 15.65
  • Mean altitude loss in last 24 hours 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 28872
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.